OpenStudy (anonymous):

will medal!!!!!!!!!!!!! 3. A model rocket is launched from the ground with an initial velocity of 352 ft/sec. d. How long will it take the rocket to reach its maximum height? Show all work in the space provided.

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hi n thnx for hlping

OpenStudy (misty1212):

same as last one \[-\frac{b}{2a}=-\frac{352}{2\times (-16)}\]

OpenStudy (anonymous):

e. Assume the model rocket’s parachute failed to deploy and the rocket fell back to the ground. How long would it take the rocket to return to Earth from the time it was launched? Show all work in the space provided.